Hi, I have the solutions to a problem but I don't know if they are correct. The problem is x^3-x=0. The solutions I got are x=-3/2 and-1/2. If they are not correct, can you show me how to do them? Thanks.
Answer by fractalier(2101)

Okay, from
x^3 - x = 0
we factor out x first and then factor what is left...
x(x^2 - 1) = 0
x(x + 1)(x - 1) = 0
so that
x = 0, x = 1, and x = -1